OpenStudy (anonymous):

When you're evaluating sec(arcsin 2/5), how would you know that it's in the first quadrant?

OpenStudy (anonymous):

or does it even matter. ;c

OpenStudy (anonymous):

It does not matter. In this question Drawing this would be helpful  To find arcsin(2/5) we need to find a number x such that sin(x)=2/5 So we draw the triangle with an angle x and 2 as our opposite side and 5 as hyp since sine= opp/hyp. So since arcsin2/5 =x all we need to find is sec x which is the same as 1/cosx I'm sure you can find cos given that diagram.

OpenStudy (anonymous):

Yees, tysm C: